# AudioBench: A Universal Benchmark for Audio Large Language Models

### v2

@@ 3.3.2 Audio Scene Understanding
Audio Captioning (AC). This task involves generating captions (descriptions) for an audio clip. To assess this task, we incorporate two popular datasets: WavCaps and AudioCaps. The examples can be found at Table 9.
@@ 3.3.3 Voice Understanding
Emotion Recognition (ER). Emotion is a critical paralinguistic element that can be conveyed through human speech or non-speech content El Ayadi et al. (2011). It plays a vital role in making communications interpretable for listeners. Three datasets are incorporated and the examples can be found at Table 7.
Accent Recognition (AR). Accent, a significant paralinguistic feature, has been overlooked in previous evaluations. We have included metadata from the VoxCeleb1 dataset and tasked the model with predicting the speaker’s likely origin based on their accent. Examples can be found at Table 10.
Gender Recognition (GR). Gender could be recognized based on vocal characteristics. This task tests the model’s sensitivity to gender-specific acoustic patterns. We have prepared two datasets and the examples can be found at Table 11.
@@ 4 Results and Analysis
Models. For comparative analysis, we selected three representative AudioLLMs: SALMONN Tang et al. (2024), Qwen-Audio Chu et al. (2023), and WavLLM Hu et al. (2024a). These models are among the most capable multi-task speech-text models available, designed to simultaneously manage a variety of tasks. Additionally, we include a cascade model named Whisper+Llama3, which processes data in a pipeline manner. First, transcriptions are extracted using the Whisper-larg-v3 Huang and Tsai (2023) model, and then these transcriptions along with user queries are input into the Llama-3-8B-Instuct model to generate responses. While this model lacks the ability to comprehend rich audio content, relying solely on transcriptions for context, it demonstrates strong performance in certain speech-intensive tasks and serves as a robust baseline.
Robustness to Diverse Queries. In evaluation, we found that AudioLLMs could behave differently to different prompt templates. Even though this also applies to text-based LLMs, it is even more severe in AudioLLMs as the two modalities are not aligned in the first stage and aggressive alignment would hurt the generalizability to different prompts. Hu et al. (2024a) also spot such issues but only limited experimental studies are presented. To study such effect, we did a more thorough evaluation in this case with 2 models (SALMONN and Qwen-Audio) on 3 ASR datasets (LirbriSpeech-Clean, CommonVoice, Tedium3).
Figure 2 illustrates the results. We can witness that SALMONN can effectively perform ASR tasks and achieve performance comparable to the Qwen-Audio model on the LibriSpeech-Clean and Tedium3 tasks, provided the right prompts are used. Specifically, prompt 3, "Decode the audio and give me the written transcripts," triggers the SALMONN model to conduct phoneme recognition on the LibriSpeech and Tedium3 datasets. For the CommonVoice dataset, SALMONN tends to perform speech translation on a significant number of samples, which adversely affects its performance on the Word Error Rate (WER) metric. This suggests that SALMONN may be overly tuned to speech features (tokens) and not sufficiently responsive to the prompts. In contrast, Qwen-Audio’s performance remains stable across all three prompt templates. Consequently, to minimize variability and thoroughly test model robustness, we incorporate at least 20 diverse prompt templates into our evaluation framework for the tasks without diverse prompts originally.
Model-as-Judge Comparison. In many datasets, we use an open-ended generation style, which is considered the most natural method of interaction. However, this style introduces challenges for evaluation, as traditional metrics are not easily adaptable. To address this, we employ a model-as-judge approach to assess model responses against reference answers and questions. Yet, there are ongoing challenges in selecting appropriate models to ensure both accessibility and robustness.
The most frequently used judge model is GPT-4. However, GPT-4 is closed-source, subject to ongoing updates, and often incurs costs that may not be feasible for the wider community. To explore alternatives, we examined 3 open-source models: Llama-3-8B-Instruct, Llama-3-70B-Instruct, and Prometheus-2 Kim et al. (2024). Prometheus-2 is specifically fine-tuned for response grading, while Llama-3-Instruct is a general-purpose open-source model.
To identify the most suitable open-source model for evaluation tasks, we use all samples from the following three datasets: CN-College-Listen, Clotho-AQA, and VoxCeleb1-Accent, to evaluate their correlation with GPT-4. The three datasets cover speech understanding, audio scene understanding, and voice understanding. Initially, we get the model outputs from the SALMONN model. These model outputs, along with golden answers and questions, are then fed into the four judgment models to obtain judgment scores. We subsequently calculated the Spearman’s rank correlation to compare the effectiveness of the different judgment models. The evaluation template can be found in Table 3. The results are detailed in Figure 3.
The Llama-3-70B-Instruct model exhibits the highest correlation with GPT-4-as-a-judge (gp4-turbo-0409). The correlation scores exceeding 0.85 across all three datasets indicate a "very strong correlation" with GPT-4. We attribute this to Llama-3-70B-Instruct’s robust generalizability. In contrast, Prometheus-2, though fine-tuned from the Mistral model and further adapted for scoring purposes, does not effectively compensate for the limitations of its base model. Therefore, in subsequent experiments, we will use Llama-3-70B-Instruct as the judging model (score 0 to 5), as it provides transparency and offers ease of adaptability. It is important to note that using a model-as-judge is not an ideal solution for grading open-ended generations and has several shortcomings. The issue of accurately grading free-form generations remains an unresolved challenge.
@@
Main Results Table 2 detailed the overall results across 26 datasets for 4 models, revealing that no single model consistently outperforms the others on all tasks. The SALMONN model is sensitive to varying instructions for ASR tasks, while Qwen-Audio and WavLLM demonstrate robust ASR capabilities. However, these AudioLLMs struggle with long-form ASR tasks without adequate context. Additionally, the training and pre-training datasets for AudioLLMs are significantly smaller than those used for other acoustic models like Whisper, which limits their generalizability to unnatural truncation.
@@
For speech-intensive tasks such as SQA and SI, the cascade model Whisper+Llama3 exhibits superior performance. This effectiveness stems from the Whisper model’s robust recognition capabilities and Llama’s strong reasoning abilities, which together efficiently handle queries primarily contained in the verbal content. In contrast, the modality fusion process in AudioLLMs may distort speech content, highlighting an area for potential improvement.
@@ Appendix A Appendix
Model-as-Judge Templates. Table 3 the scoring templates used for the model-as-judge approach. The scoring range is from 0 to 5, except for Prometheus2, where the range is from 1 to 5.
@@
Dataset Context Instruction (Sample) Answer LibriSpeech-Clean [Uncaptioned image] NO I WASN’T THINKING OF THAT I need the spoken content written down, please. NO I WASN’T THINKING OF THAT LibriSpeech-Other [Uncaptioned image] THE HISTORY OF THE HOUSE IS PLAIN NOW I’m asking for a transcription of the spoken words into text. THE HISTORY OF THE HOUSE IS PLAIN NOW CommonVoice [Uncaptioned image] This is Jon Davis. Could you write out what was said? This is Jon Davis. PeoplesSpeech [Uncaptioned image] that’s where you have a lot of windows in the south no actually that’s passive solar and passive solar is something that was… Could you transcribe what you hear? that’s where you have a lot of windows in the south no actually that’s passive solar and passive solar is something that was… GigaSpeech [Uncaptioned image] I DO THINK THAT FAST FASHION HAS KIND OF ALLOWED EVERYONE TO HAVE ACCESS TO AESTHETICALLY THOUGHTFUL CLOTHING. Capture the speech in written format, please. I DO THINK THAT FAST FASHION HAS KIND OF ALLOWED EVERYONE TO HAVE ACCESS TO AESTHETICALLY THOUGHTFUL CLOTHING <PERIOD> Earning-21 [Uncaptioned image] TOKEN GOOD MORNING EVERYONE AND WELCOME TO THE NEXTERA ENERGY INC AND NEXTERA ENERGY PARTNERS… Please document the spoken language. TOKEN GOOD MORNING EVERYONE AND WELCOME TO THE NEXTERA ENERGY INC AND NEXTERA ENERGY PARTNERS… Earning-22 [Uncaptioned image] TOKEN GOOD DAY LADIES AND GENTLEMEN AND WELCOME TO THE CD PROJEKT GROUP FINANCIAL RESULTS FOR H1 2021 CONFERENCE CALL TODAY’S CALL … Document the conversation by writing it down. TOKEN GOOD DAY LADIES AND GENTLEMEN AND WELCOME TO THE CD PROJEKT GROUP FINANCIAL RESULTS FOR H1 2021 CONFERENCE CALL TODAY’S CALL … Tedium3 [Uncaptioned image] one day los angeles times columnist steve lopez was walking along the streets … Please transcribe the speech. one day los angeles times columnist steve lopez was walking along the streets … Tedium3-Longform [Uncaptioned image] i ’d like to share with you a discovery that i made a few months ago while writing an article for italian wired i always keep my thesaurus handy whenever … Please transcribe the speech. i ’d like to share with you a discovery that i made a few months ago while writing an article for italian wired i always keep my thesaurus handy whenever …
@@
Table 4: Examples from Automatic Speech Recognition (ASR) datasets
@@
Dataset Context Instruction Answer CN-College-Exam-Listening [Uncaptioned image] F: Excuse me, this is the address, where do I find it? M: All right, you need a street map, here is one, and I will show you where it is. What does the woman want to do? Find a place. SLUE-P2-SQA5 [Uncaptioned image] 1 Climate Stephens City is located in the humid subtropical climate zone ( Köppen climate classification : Cfa ) … Which regions have temperate climates? mid-atlantic DREAM-TTS [Uncaptioned image] F: The movie next Tuesday has been canceled due to lack of interest. M: what do you mean? F: Well by last night only a few tickets been sold. What can we conclude about the movie? The movie will not be shown. Public-SG-SpeechQA [Uncaptioned image] Today, speaking to a roomful of economists, I am inclined to confine myself to talk about markets and price, dollars … How can economics help solve complex healthcare challenges, as mentioned by the speaker? Economics can help solve complex healthcare …
@@
Dataset Context Instruction (Sample) Answer IEMOCAP-Emotion [Uncaptioned image] Thank you. How would you interpret the speaker’s emotional state from their speech? From the speaker’s speech, it seems they are in a sad state. MELD-Sentiment [Uncaptioned image] Yeah, I’m not in that. How would you interpret the speaker’s sentiment state from their speech? From the speaker’s speech, it seems they are in a neutral sentiment state. MELD-Emotion [Uncaptioned image] Yeah, I’m not in that. What do you think the speaker is feeling based on their speech patterns (e.g., pace, pitch, volume)? Based on the speaker’s speech patterns, it seems like they are feeling neutral.
@@
Dataset Context Instruction (Sample) Answer VoxCeleb-Accent [Uncaptioned image] because I was trying to get away from her, but where did you go? what did you do? From which country does the speaker hail? From the audio, I guess the speaker is from USA.
@@
Dataset Context Instruction (Sample) Answer VoxCeleb-Gender [Uncaptioned image] because I was trying to get away from her, but where did you go? what did you do? Is there information available on the speaker’s gender? The speaker is a female. IEMOCAP-Gender [Uncaptioned image] For God’s sake, Augie. It’s- Grow up, we’re not going to see the grunion. Based on what you heard, can you identify whether the speaker is male or female? Based on the auditory cues, it sounds like the speaker is a male.
@@
Figure 5: OpenHermes-Audio Labeling Platform.
Figure 6: ALPACA-Audio Labeling Platform.
Figure 7: Labeling Platform for AudioCaps QA and WavCasp QA.
Dataset Name Metrics AudioLLMs Whisper+Llama3 SALMONN Qwen-Audio WavLLM Speech Understanding LibriSpeech-Clean WER 45.28 2.20 2.05 1.83 LibriSpeech-Other 33.76 4.40 4.88 3.71 CommonVoice-15 44.44 11.30 14.84 9.89 PeoplesSpeech 35.05 30.59 37.51 14.55 GigaSpeech 14.42 13.18 16.02 9.52 Tedlium3 13.21 4.28 5.43 3.81 Tedlium3-Longform 25.88 34.16 46.51 4.94 Earning-21 31.81 39.39 65.64 11.77 Earning-22 35.87 54.37 67.22 15.60 CN-College-Listen M.J.8B 1.57 1.63 1.95 2.32 SLUE-P2-SQA5 3.09 3.03 3.08 3.26 DREAM-TTS 1.46 1.35 1.86 2.23 Public-SG-SpeechQA 2.18 2.42 2.18 2.51 OpenHermes-Audio 2.15 2.12 1.72 2.34 ALPACA-Audio 1.72 2.19 1.71 2.39 Audio Scene Understanding Clotho-AQA M.J.8B 2.08 2.26 1.76 1.28 WavCaps-QA 1.90 1.58 1.03 0.95 AudioCaps-QA 1.94 1.93 1.06 1.05 WavCaps 1.18 1.16 0.51 0.65 AudioCaps 1.47 1.39 0.49 0.56 WavCaps METEOR 21.36 21.75 9.78 14.52 AudioCaps 22.79 22.91 6.70 7.15 Voice Understanding IEMOCAP-Emotion M.J.8B 1.59 1.00 1.71 1.44 MELD-Sentiment 1.51 1.21 1.30 1.93 MELD-Emotion 1.16 1.02 1.04 1.47 VoxCeleb1-Accent 1.52 1.59 1.38 0.95 VoxCeleb1-Gender 2.58 2.74 2.41 1.67 IEMOCAP-Gender 1.35 1.95 1.54 1.56
Table 12: Main evaluation results of various AudioLLMs and one cascade model. M.J.8B refers to Model-as-Judge with Llama-3-8B-Instruct model (Score is from 0-lowest to 5-highest). METEOR scores are common metrics for audio captioning Banerjee and Lavie (2005). WER is the word error rate for ASR tasks. When outperforming end-to-end models, the Whisper+Llama3 model is marked with underline.

### v3

@@ 3.3.2 Audio Scene Understanding
Audio Captioning (AC). This task involves generating captions (descriptions) for an audio clip. We incorporate two popular datasets to assess this task: WavCaps and AudioCaps. The examples can be found in Table 9. Banerjee and Lavie (2005)
@@ 3.3.3 Voice Understanding
Emotion Recognition (ER). Emotion is a critical paralinguistic element that can be conveyed through human speech or non-speech content El Ayadi et al. (2011). It plays a vital role in making communications interpretable for listeners. Three datasets are incorporated and the examples can be found in Table 7.
Accent Recognition (AR). Accent, a significant paralinguistic feature, has been overlooked in previous evaluations. We have included metadata from the VoxCeleb1 dataset and tasked the model with predicting the speaker’s likely origin based on their accent. Examples can be found in Table 10.
Gender Recognition (GR). Gender could be recognized based on vocal characteristics. This task tests the model’s sensitivity to gender-specific acoustic patterns. We have prepared two datasets and the examples can be found in Table 11.
@@ 4.1 Main Results
For comparative analysis, we selected four representative AudioLLMs: SALMONN Tang et al. (2024), Qwen-Audio Chu et al. (2023), WavLLM Hu et al. (2024a) and Qwen2-Audio Chu et al. (2024). These models are among the most capable multi-task speech-text models available, designed to manage various tasks simultaneously. Additionally, we include a cascade model named Whisper+Llama3, which processes data in a pipeline manner. First, transcriptions are extracted using the Whisper-larg-v3 Huang and Tsai (2023) model, and then these transcriptions along with user queries are input into the Llama-3-8B-Instuct model to generate responses. While this model cannot comprehend rich audio content, relying solely on transcriptions for context, it demonstrates strong performance in speech-intensive tasks and serves as a robust baseline.
@@
Table 2 detailed the overall results across 26 datasets for 5 models, revealing that no single model consistently outperforms the others on all tasks. Specifically, SALMONN model is sensitive to varying instructions for ASR tasks, which indicates that it is overfitting to certain audio features and ignores proper instructions. Qwen1/2-Audio and WavLLM demonstrate robust ASR capabilities. However, all AudioLLMs struggle with long-form ASR tasks. This could be because the models are mainly fine-tuned with limited audio length which makes them hard to generalize to arbitrary context lengths. Additionally, the training and pre-training datasets for AudioLLMs are generally smaller than those used for other task-specific models like Whisper, which limits their generalizability to unnatural truncation as well.
@@
For speech-intensive tasks such as SQA and SI, the cascade model Whisper+Llama3 exhibits superior performance. This effectiveness stems from the Whisper model’s robust recognition capabilities and Llama’s strong reasoning abilities, which together efficiently handle queries primarily contained in the verbal content. In contrast, the modality fusion process in AudioLLMs may distort speech content, highlighting an area for future improvement.
@@
Figure 3: Correlation results for three benchmark datasets across three variations of the model-as-judge approach were analyzed using Spearman’s rank correlation. The findings show that Llama-3-70B-Instruct has a very strong correlation with GPT-4, when serving as the judge.
4.2 Robustness Queries
In evaluation, we found that AudioLLMs could behave differently to different prompt templates. Even though this also applies to text-based LLMs, it is even more severe in AudioLLMs as the two modalities are not aligned in the first stage and aggressive alignment would hurt the generalizability of different prompts. Hu et al. (2024a) also spot such issues but only limited experimental studies are presented. Therefore, we did a more thorough experimental evaluation in this case with 2 models (SALMONN and Qwen-Audio) on 3 ASR datasets (LirbriSpeech-Clean, CommonVoice, Tedium3).
Figure 2 illustrates the results. We can witness that SALMONN can effectively perform ASR tasks and achieve performance comparable to the Qwen-Audio model on the LibriSpeech-Clean and Tedium3 tasks, provided the right prompts are used. However, prompt 3, "Decode the audio and give me the written transcripts," triggers the SALMONN model to conduct phoneme recognition on the LibriSpeech and Tedium3 datasets. For the CommonVoice dataset, SALMONN tends to perform speech translation on a significant number of samples, which adversely affects its performance on the Word Error Rate (WER) metric. This suggests that SALMONN may be overly tuned to speech features (tokens) and not sufficiently responsive to the prompts. In contrast, Qwen-Audio’s performance remains stable across all three prompt templates. Consequently, to minimize variability and thoroughly test model robustness, we incorporate at least 20 diverse prompt templates into our evaluation framework for the tasks without diverse prompts originally.
4.3 Model-as-Judge Comparison
In many datasets, we use an open-ended generation style, which is considered the most natural method of interaction. However, this style introduces challenges for evaluation, as traditional metrics are not easily adaptable. To address this, we employ a model-as-judge approach to assess model responses against reference answers and questions. Yet, there are ongoing challenges in selecting appropriate models to ensure both accessibility and robustness.
The most frequently-used judgement model is GPT-4. However, GPT-4 is closed-source, subject to ongoing updates, and often incurs costs that may not be feasible for the wider community. To explore alternatives, we examined 3 open-source models: Llama-3-8B-Instruct, Llama-3-70B-Instruct, and Prometheus-2 Kim et al. (2024). Prometheus-2 is specifically fine-tuned for response grading, while Llama-3-Instruct is a general-purpose open-source model.
To identify the most suitable open-source model for evaluation tasks, we use all samples from the following three datasets: CN-College-Listen, Clotho-AQA, and VoxCeleb1-Accent, to evaluate their replaceability of GPT-4. The three datasets cover speech understanding, audio scene understanding, and voice understanding. Initially, we get the model outputs from the SALMONN model. These model outputs, along with golden answers and questions, are then fed into the four judgment models to obtain judgment scores. We subsequently calculated the Spearman’s rank correlation to compare the effectiveness of the different judgment models. The evaluation template can be found in Table 3. The results are detailed in Figure 3.
As a result, the Llama-3-70B-Instruct model exhibits the highest correlation with GPT-4-as-a-judge (gp4-turbo-0409). The correlation scores exceeding 0.85 across all three datasets indicate a "very strong correlation" with GPT-4. We attribute this to Llama-3-70B-Instruct’s robust generalizability. In contrast, Prometheus-2, though fine-tuned from the Mistral model and further adapted for scoring purposes, does not effectively compensate for the limitations of its base model. Therefore, in experiments, we will use Llama-3-70B-Instruct as the judging model (scale from 0-5 or 0-1 followed by rescaling to 100-point scale for easier interpretation), as it provides transparency and offers ease of adaptability. It is important to note that using a model-as-judge is not an ultimate solution. The issue of accurately grading free-form generations remains an unresolved challenge in NLP.
@@ A.1 Discussion on Dataset
A.2 Model-as-Judge Template
Table 3 the scoring templates used for the model-as-judge approach. The scoring range is from 0 to 5, except for Prometheus2, where the range is from 1 to 5. For some tasks that require binary judgment, the score will be either 0-incorrect or 1-correct. In practice, we rescaled all scores to a 100-point scale for easier comparison. For some tasks,
A.3 Comparison with AIR-Bench
AIR-Bench is one concurrent work with AudioBench Yang et al. (2024a). Even though both focus on evaluating audio-based instructions with LLMs. The selected datasets and evaluation preferences are different. First, the biggest difference comes from the coverage of datasets. In AudioBench, we proposed 6 new datasets to accommodate what is missing in the evaluation. Meantime, we have multiple ASR datasets, speech question answer datasets, and speech instruction datasets that are not covered by AIR-Bench. Second, the AudioBench evaluation toolkit takes care of prompt variants and considers the model’s robustness in following instructions. Third, we present a holistic study on the choice of evaluation metrics and provide stable and affordable solutions for future benchmarking purposes. Some merits can also be seen in AIR-Bench which also includes Music datasets. Nevertheless, AudioBench servers a good foundation for evaluating AudioLLMs in three aspects: speech understanding, audio scene understanding, and voice understanding, and are expandable toolkit for more suitable scenarios.
A.4 Why SUPERB and Dynamic-SUPERB is not Ideal
For instance, a pair ($Audio$, $Query$) is presented to the model to elicit a ($Model\_Answer$), which is then compared with a reference answer for relevant tasks. This approach does not align well with previous speech benchmarks like SUPERB Yang et al. (2024b) and Dynamic-SUPERB Huang et al. (2024a). The design for SUPERB is for the evaluation of a self-supervised speech encoder where a supervised fine-tuning step is normally conducted for the evaluation. Dynamic-SUPERB can measure the instruction following tasks in a zero-shot fashion. However, it is an open collection with crowd-sourced contributions that lacks a specific focus for AudioLLMs as also discussed in Yang et al. (2024a). Therefore, AudioBench is introduced to focus primarily on evaluating AudioLLMs, offering new datasets and evaluation pipelines to address the existing gaps.
@@ A.4 Why SUPERB and Dynamic-SUPERB is not Ideal
Dataset Context Instruction (Example) Answer LibriSpeech-Clean [Uncaptioned image] No, I wasn’t thinking of that. Turn the speech input into a text transcription. No, I wasn’t thinking of that. LibriSpeech-Other [Uncaptioned image] The history of the house is plain now. Decode the audio and give me the written transcription. The history of the house is plain now. CommonVoice [Uncaptioned image] This is Jon Davis. Process the audio speech and provide the text output. This is Jon Davis. PeoplesSpeech [Uncaptioned image] that’s where you have a lot of windows in … Convert the audio speech into a text transcript. that’s where you have a lot of windows in … GigaSpeech [Uncaptioned image] I believe that fast fashion has made it possible for everyone to have access to aesthetically thoughtful clothing. Transform the speech into a text document. I believe that fast fashion has made it possible for everyone to have access to aesthetically thoughtful clothing. Earning-21 [Uncaptioned image] Good morning, everyone, and welcome to the NextEra Energy Inc. and NextEra Energy Partners… Convert the audio speech into a text transcript. Good morning, everyone, and welcome to the NextEra Energy Inc. and NextEra Energy Partners… Earning-22 [Uncaptioned image] Good day, ladies and gentlemen, and welcome to the CD Projekt Group financial results for H1 2021 conference call. Today’s call… Transform the speech into a text document. Good day, ladies and gentlemen, and welcome to the CD Projekt Group financial results for H1 2021 conference call. Today’s call… Tedium3 [Uncaptioned image] One day, Los Angeles Times columnist Steve Lopez was walking along the streets… Turn the speech input into a text transcription. One day, Los Angeles Times columnist Steve Lopez was walking along the streets… Tedium3-Longform [Uncaptioned image] I’d like to share with you a discovery I made a few months ago while writing an article for Italian Wired. I always keep my thesaurus handy whenever… Process the audio speech and provide the text output. I’d like to share with you a discovery I made a few months ago while writing an article for Italian Wired. I always keep my thesaurus handy whenever…
@@
Table 4: Examples from Automatic Speech Recognition (ASR) datasets. Capitalization normalized for display purposes.
@@
Dataset Context Instruction Answer CN-College-Exam-Listening [Uncaptioned image] F: Excuse me, this is the address, where do I find it? M: All right, you need a street map, here is one, and I will show you where it is. Question: What does the woman want to do? Choices: (A) Find a place. (B) Buy a map. (C) Get an address. (A) Find a place. SLUE-P2-SQA5 [Uncaptioned image] 1 Climate Stephens City is located in the humid subtropical climate zone ( Köppen climate classification : Cfa ) … Which regions have temperate climates? mid-atlantic DREAM-TTS [Uncaptioned image] F: The movie next Tuesday has been canceled due to a lack of interest. M: what do you mean? F: Well by last night only a few tickets been sold. Question: What can we conclude about the movie? Choices: (A) They want to buy the tickets for the movie. (B) The tickets for the movie were sold. (C) The movie will not be shown. (C) The movie will not be shown. Public-SG-SpeechQA [Uncaptioned image] Today, speaking to a roomful of economists, I am inclined to confine myself to talk about markets and price, dollars … How can economics help solve complex healthcare challenges, as mentioned by the speaker? Economics can help solve complex healthcare …
@@
Dataset Context Instruction (Sample) Answer IEMOCAP-Emotion [Uncaptioned image] Thank you. Can you interpret the emotions in the speaker’s speech (frustration, anger, excited, neutral, happiness, surprise, sad)? From the speaker’s speech, it seems they are in a sad state. MELD-Sentiment [Uncaptioned image] Yeah, I’m not in that. What sentiment signals can you hear in the speaker’s speech (neutral, positive, negative)? From the speaker’s speech, it seems they are in a neutral sentiment state. MELD-Emotion [Uncaptioned image] Yeah, I’m not in that. How does the speaker’s speech reflect their emotional state (neutral, joy, disgust, sadness, surprise, anger, fear)? Based on the speaker’s speech patterns, it seems like they are feeling neutral.
@@
Dataset Context Instruction (Sample) Answer VoxCeleb-Accent [Uncaptioned image] because I was trying to get away from her, but where did you go? what did you do? Based on their accent, where is the speaker most likely from? From the audio, I guess the speaker is from USA.
@@
Dataset Context Instruction (Sample) Answer VoxCeleb-Gender [Uncaptioned image] because I was trying to get away from her, but where did you go? what did you do? Can you determine the gender of the speaker from the audio (Male or Female)? The speaker is a female. IEMOCAP-Gender [Uncaptioned image] For God’s sake, Augie. It’s- Grow up, we’re not going to see the grunion. From the audio, can you determine the speaker’s gender (Male or Female)? Based on the auditory cues, it sounds like the speaker is a male.
@@
Figure 5: Our Labeling Platform for OpenHermes-Audio.
Figure 6: Our Labeling Platform for ALPACA-Audio.
Figure 7: Our Labeling Platform for AudioCaps QA and WavCasp QA.